A car is stopped at a traffic light. It then travels along a straight road such



that its distance from the light is given by:



π‘₯(𝑑) = 2.4𝑑



2 βˆ’ 0.120𝑑



3

1
Expert's answer
2022-03-03T14:26:22-0500

"x=2.4t^2-0.120t+3"

Differentiate with respect to time

"v=\frac{dx}{dt}"

"v=\frac{d(2.4t^2-.120t+3)}{dt}=4.8t-0.120"

"v=0"

"4.8t-0.120=0"

"t=\frac{0.120}{4.8}=0.025sec"